# Multilayer network science: theory, methods, and applications

### v1

@@ II.1.4 Multilayer metrics
Some metrics and concepts do not have counterparts in standard single-layer networks. The characterization of multilayer networks through these new metrics has unveiled an array of phenomena hitherto unobserved on single-layered systems, thus manifesting the singularity and usefulness of the multilayer description. For instance, the participation coefficient measures how evenly a node’s links are spread across the different layers. A node with high multi-degree but low participation coefficient is a hub - a highly connected node - in only one or a few layers, while a node with high participation coefficient is genuinely active across the entire system. This distinction is crucial for understanding a node’s role in processes that span multiple layers battiston2014structural.
Another network-level metric with no single-layer equivalent is the mutually largest connected component (MLCC). This is the largest group of nodes where every node can reach every other node through paths existing within each and every layer buldyrev_2010. The size of the MLCC is a powerful indicator of a system’s integrated resilience artime2024robustness. In many multilayer systems, reducing network connectivity (e.g., by removing nodes or links) can cause the MLCC to collapse abruptly in a discontinuous phase transition, contrasting sharply with the smoother, continuous transitions typically seen in single-layer networks gao2012networks; radicchi2013abrupt
A fundamental question in multilayer network science is: “What new information does the layered structure provide compared to a simple, aggregated network?” To answer this, we need metrics that quantify the relationships between layers. One approach is to measure layer similarity or redundancy. This can be done by calculating the interlayer mutual information or other correlation measures between the adjacency matrices of different layers. If layers are highly similar, the system is redundant, and an aggregated view might suffice. If they are diverse, the multilayer representation is essential, as different layers capture distinct relational patterns DeDomenico2015ranking; Schieber2017; ghavasieh2020enhancing.
Furthermore, multilayer networks introduce new types of structural correlations. Beyond intralayer patterns like degree assortativity (whether high-degree nodes connect to other high-degree nodes pastor01b; newman02), we can measure interlayer correlations. For example, are nodes with a high degree in one layer also likely to have a high degree in another? The presence or absence of such correlations significantly impacts the behavior of dynamical processes, from spreading to synchronization, and has been a major focus of research pamfil2019relating; nicosia2015measuring; wu2020correlated; nicosia2013growing; bianconi2013statistical; battiston2014structural; menichetti2014weighted; lee2012correlated.
@@ II.2.1 Community detection
How shall we define communities in multilayer networks? The answer is context-dependent. In a network of customers layered by product categories, a community might be a group of people with consistently similar purchasing habits across multiple layers (a cross-market segment). In a biomedical context, coupling a gene co-expression network with a protein-protein interaction network, a community could reveal groups of genes that are co-expressed and whose encoded proteins physically interact. The search for such meaningful groups in multilayer networks has become a vibrant research area, with several reviews already charting the landscape kim15; magnani21; huang21. Here, we summarize the main strategies and highlight key challenges and future directions.
@@
The first strategy, flattening, collapses the multilayer network into a single aggregated graph containing all nodes and links from all layers berlingerio11; kim16. Any traditional algorithm working on single-layer networks can then be applied. While computationally efficient, this approach discards all information about the distinct layers and their relationships.
Layer-by-layer algorithms first detect communities on each layer independently and then aggregate the results. These are further divided into three classes: frequent subgraph mining zeng06; berlingerio13, feature integration dong13; nie18, and partition integration cantini15; tagarelli17. Frequent subgraph mining is based on the idea that the nodes of a multilayer community are assigned to the same group in different layers. Thus, it relies on identifying groups of nodes that are consistently clustered together across many layers. Feature integration relies on the principle that communities are made of nodes that are similar to each other, something typically measured by looking at the eigenvectors of the layers. Communities are found by grouping nodes according to their features via, e.g., k-means clustering forgy65; macqueen67. Partition integration directly combines the divisions found on the individual layers. For instance, one could derive a consensus partition, which is the division that is closest to all the input ones, on average strehl02.
Figure 3: Multilayer communities. Communities are indicated by the dashed contours. (a) Intralayer communities group state nodes on the same layer. (b) Members of cross-layer communities can be on both the same and on different layers. (c) Pillar communities group all the counterparts of the nodes in other layers in the same community.
@@
Many of these methods are a direct extension of single-layer algorithms to the multilayer case. Modularity maximization was generalized by defining a multilayer null model based on random walks, with intralinks and interlinks playing different roles mucha10. Local exploration methods, like seed set expansion interdonato17, label propagation boutemine17 or the clique percolation method afsarmanesh18, as well as random walk-based algorithms such as Infomap dedomenico15 and Walktrap kuncheva15, have also been adapted. Furthermore, statistical inference methods, primarily based on the stochastic block model (SBM), offer a principled framework to detect communities, infer relationships between layers, and even test whether an observed single-layer network is better explained as an aggregation of hidden layers peixoto15; debacco17.
Despite this progress, significant challenges remain. First, a persistent issue is the rigorous assessment of statistical significance. Without it, one cannot distinguish meaningful communities from patterns arising by chance. Second, the increasing availability of node and edge metadata is often used only for validating results, a procedure known to be unreliable hric14; peel17. Instead, this information could be directly integrated into the detection process to find more accurate and relevant communities, a task for which SBM-based approaches are particularly well-suited hric16; newman16. Finally, the vast majority of methods are designed for multiplex networks. There is a clear need for algorithms that can handle more general multilayer networks (often called heterogeneous networks in computer science) with different node sets and relationships in each layer, a promising direction for future research sun13.
@@ II.2.2 Dynamics and Processes
A paradigmatic case in which multiplexity alters the behaviour of a well-understood process is that of reaction-diffusion systems. In multiplex networks, reaction and diffusion terms can be assigned to different layers or coupled across them. For instance, in Kouvaris2015, activator and inhibitor species are placed on separate layers, where each species diffuses only within its own layer but reacts across layers. Remarkably, the presence of interlayer coupling allows for Turing-like pattern formation even when both species have identical mobility rates—something impossible in single-layer networks. This form of topology-driven instability, further explored in Asllani2014, highlights how multiplexity can give rise to heterogeneous and robust spatial patterns under conditions where the dynamics of isolated layers remain homogeneous.
@@
In multiplex systems, one well-studied example of interacting dynamics is the coupling between epidemic spreading and awareness diffusion Granell2013Sep. In this model, the disease propagates through physical contact on one layer, while information about the disease (awareness) spreads via virtual communication on a separate layer. The two processes mutually inhibit each other, giving rise to a point at which the presence of awareness fundamentally alters both the epidemic threshold and the eventual prevalence.
@@
Multiplex networks also provide a natural setting for studying feedback between processes on different layers. In Nicosia2017, a model is introduced in which neural synchronization on one layer modulates resource allocation on another, and vice versa. This mutual feedback leads to the spontaneous emergence of explosive synchronization and highly heterogeneous resource distributions.
@@
Multiplex networks often feature layers with different topologies or dynamics, leading to hybrid systems. In Hizanidis2016, a modular neural network is modelled with electrical synapses within communities and chemical synapses across them, each type represented as a different layer. The interplay between these types of couplings gives rise to chimera-like states, where coherent and incoherent activity coexist across modules. Such hybrid multiplex models are relevant to understanding how local synchrony coexists with global disorder in brain dynamics.
@@
A similar approach is taken in Sadilek2015, where a two-layer Kuramoto model is derived from physiologically motivated equations for cortical dynamics. The model captures transitions between resting and background brain activity, successfully reproducing empirical features that were inaccessible to simpler models. Again, multiplexity allows for distinct coupling schemes across layers (e.g., electrical vs chemical, or fast vs slow processes) that shape the global phase diagram.
@@
Another work examined how two layers not directly connected can synchronize through the mediation of a central relay layer Leyva2018. Results showed that synchronization thresholds are reduced in the multiplex setting and that low-degree nodes play a crucial role, while hubs can be “demultiplexed” without disrupting global coherence. More recently, a three-layer multiplex of non-linear oscillators, in which a central drive layer unidirectionally couples to two response layers, has been investigated Vadakkan2025. Such directional coupling induces synchronized oscillations in the response layers, even with amplification relative to the drive, and varying interlayer strength or introducing time-scale mismatches enables control over synchronization patterns, amplitudes, and dynamical regimes such as quasi-periodicity and generalized synchronization.
@@ II.3.1 Temporal Networks
Complex systems, from social groups to interacting neurons, are fundamentally dynamic. While classic network science often represents these systems with static graphs, this is an aggregation that loses all temporal information, such as the sequence, duration, or concurrency of interactions. The recent availability of time-resolved data and computational power has propelled the study of temporal networks, revealing phenomena invisible in a static view holme2012temporal; holme2015modern. A temporal network can be represented as a sequence of static snapshots, which naturally fits the multilayer network framework where each layer is a snapshot in time. Besides, it is also possible to have networks that are temporal and multilayer at the same time, in which case the temporal correlations between layers can significantly affect the dynamics of processes unfolding on multilayer networks starnini_effects_2017.
@@
While early theoretical models added temporality by dynamically changing links, particularly in an adaptive fashion gross2008adaptive, the strong expansion of the field has been mostly data-driven holme2012temporal; holme2015modern. The availability of high-resolution temporal data, such as the face-to-face contact network in a museum exhibition depicted in Figure 4 and collected by the SocioPatterns collaboration using wearable sensors cattuto2010dynamics, provided a strong empirical foundation for the field. This data made it clear that static representations not only hide temporal details but can be misleading, as the properties of a networked system can vary dramatically from one moment to the next masuda2016guide; saramaki2015seconds.
@@
Figure 4: Temporal networks of face-to-face interactions of people on different days of museum visits. In each day, time evolution is represented as a color code. The data has been recorded by detecting physical proximity using wearable sensors developed by the SocioPatterns collaboration isella2011what; vandenbroeck2012making.
@@
This data-centric view is what led researchers to analyse temporal networks as a series of snapshots. This approach revealed that while aggregated properties might converge to the static ones krings2012effects, local measures like a node’s centrality could fluctuate significantly, with nodes shifting from central to peripheral roles over time braha2006centrality; braha2009time. However, a true paradigm shift occurred when the focus moved from fluctuating static properties to inherently temporal concepts. Researchers recognized that information flows along temporal paths, which are a set of time-respecting sequences of interactions kempe2000connectivity; holme2005network. A path that exists in an aggregated static network may be impossible in a temporal one if the interactions do not occur in the correct chronological order pan2011path. This insight spurred the redefinition of classical centrality measures for the temporal domain stehle2010dynamical and led to the discovery of new phenomena, such as the burstiness of human interactions, where events are clustered in time rather than occurring randomly barabasi2005origin; karsai2018bursty.
@@
But not all temporal interactions are equally informative. Patterns may appear briefly, recur, or vanish entirely. Motifs that are evident in aggregated static views of the network may never exist in real time due to improper chronological alignment holme_network_2005. To deal with this complexity, researchers have developed various techniques for filtering and pattern detection. At the local scale, temporal backbones highlight node pairs whose interaction frequencies significantly exceed what would be expected under a temporal null model kobayashi2019structured. Temporal motifs identify frequently recurring, time-ordered subgraphs kovanen2013temporal, while temporal cores generalize the static core decomposition to uncover dense substructures active over defined time windows galimberti2018mining. Similarly, the temporal rich-club coefficient measures whether highly connected nodes in the aggregate network tend to form simultaneous, persistent connections pedreschi2022temporal.
@@
In parallel, significant effort has gone to identifying the relevant timescales that govern the evolution of temporal networks. Early techniques relied on comparing successive time windows darst2016detection, but this overlooked long-range correlations. More sophisticated methods compute pairwise distances or similarities between network snapshots masuda2019detecting, enabling the clustering of recurring states and the identification of their transitions. Other approaches apply auto-correlation metrics, or Fourier transforms to extract temporal correlations and recurrent interaction patterns at different time scales lacasa2022correlations; andres2024detecting.
@@
Representing temporal networks in a simple form for analysis has also been a priority. Two lossless, static-like representations have emerged as especially useful. The supra-adjacency representation encodes node-time pairs as nodes in a multilayer structure, linking them according to the observed interactions valdano2015analytical; sato2019dyane. The event graph, in contrast, treats interactions themselves as nodes, linking temporally adjacent events involving shared participants kivela2018mapping. This latter framework is especially well-suited to model processes constrained by timing, such as epidemic spreading or transit dynamics badie2022directed. Another line of work uses graph embedding techniques to map high-dimensional temporal networks into low-dimensional vector spaces goyal18; cai18. These embeddings preserve structural or temporal similarities and support tasks such as link prediction or influence sets thongprayoon2023embedding; torricelli2020weg2vec; dall2024embedding. Embeddings also facilitate data compression and learning-based applications, as discussed further in Section II.3.3.
@@
Models of temporal networks have been developed to generate realistic synthetic networks or for establishing null benchmarks. Mechanistic models aim to reproduce observed features or test hypotheses about dynamics holme2015mechanistic, while randomized models offer a controlled way to destroy correlations and assess their significance gauvin2022randomized. These models range from memoryless activity-driven frameworks perra_activity_2012; starnini_topological_2013 to non-Markovian extensions capturing burstiness and memory effects moinet_burstiness_2015. Additional approaches include stochastic block models, priority queue models, models based on Hawkes or other dynamical processes, and memory network models.
@@
The temporal networks framework also adds another layer of complexity to dynamical processes unfolding on them, since one needs to take into account the coupling between the characteristic timescale of the dynamic process and the temporal network’s one. In epidemiology, the coupling between interactions and disease progression profoundly affects spreading outcomes with important consequences for designing effective public health interventions masuda2017introduction; machens2013infectious; koher2019contact; toroczkai2007proximity; valdano2015analytical. In social systems, temporal networks are used to model the spread of information, behaviours, and innovations, often through threshold mechanisms distinct from disease transmission unicomb2021dynamics; andres2024competition. Temporal opinion dynamics have been explored to understand consensus and polarization li2019impact; chu2023non, while synchronization phenomena in biology and engineering are modelled as coupled oscillators interacting over time-varying links ghosh2022synchronized. Applications in transportation include traffic optimization and infrastructure design salama2022temporal.

### v2

@@ II.1.4 Multilayer metrics
Some metrics and concepts do not have counterparts in standard single-layer networks. The characterisation of multilayer networks through these new metrics has unveiled an array of phenomena hitherto unobserved on single-layered systems, thus manifesting the singularity and usefulness of the multilayer description. For instance, the participation coefficient measures how evenly a node’s links are spread across the different layers. A node with high multi-degree but low participation coefficient is a hub - a highly connected node - in only one or a few layers, while a node with high participation coefficient is genuinely active across the entire system. This distinction is crucial for understanding a node’s role in processes that span multiple layers Battiston et al. (2014).
Another network-level metric with no single-layer equivalent is the mutually largest connected component (MLCC). This is the largest group of nodes where every node can reach every other node through paths existing within each and every layer Buldyrev et al. (2010). The size of the MLCC is a powerful indicator of a system’s integrated resilience Artime et al. (2024). In many multilayer systems, reducing network connectivity (e.g., by removing nodes or links) can cause the MLCC to collapse abruptly in a discontinuous phase transition, contrasting sharply with the smoother, continuous transitions typically seen in single-layer networks Gao et al. (2012); Radicchi and Arenas (2013)
A fundamental question in multilayer network science is: “What new information does the layered structure provide compared to a simple, aggregated network?” To answer this, we need metrics that quantify the relationships between layers. One approach is to measure layer similarity or redundancy. This can be done by calculating the interlayer mutual information or other correlation measures between the adjacency matrices of different layers. If layers are highly similar, the system is redundant, and an aggregated view might suffice. If they are diverse, the multilayer representation is essential, as different layers capture distinct relational patterns De Domenico et al. (2015b); Schieber et al. (2017); Ghavasieh and De Domenico (2020).
Furthermore, multilayer networks introduce new types of structural correlations. Beyond intralayer patterns like degree assortativity (whether high-degree nodes connect to other high-degree nodes Pastor-Satorras et al. (2001); Newman (2002)), we can measure interlayer correlations. For example, are nodes with a high degree in one layer also likely to have a high degree in another? The presence or absence of such correlations significantly impacts the behaviour of dynamical processes, from spreading to synchronisation, and has been a major focus of research Pamfil et al. (2019); Nicosia and Latora (2015); Wu et al. (2020); Nicosia et al. (2013a); Bianconi (2013); Battiston et al. (2014); Menichetti et al. (2014); Lee et al. (2012).
@@ II.2.1 Community detection
How shall we define communities in multilayer networks? The answer is context-dependent. In a network of customers layered by product categories, a community might be a group of people with consistently similar purchasing habits across multiple layers (a cross-market segment). In a biomedical context, coupling a gene co-expression network with a protein-protein interaction network, a community could reveal groups of genes that are co-expressed and whose encoded proteins physically interact. The search for such meaningful groups in multilayer networks has become a vibrant research area, with several reviews already charting the landscape Kim and Lee (2015); Magnani et al. (2021a); Huang et al. (2021). Here, we summarize the main strategies and highlight key challenges and future directions.
@@
The first strategy, flattening, collapses the multilayer network into a single aggregated graph containing all nodes and links from all layers Berlingerio et al. (2011); Kim et al. (2016). Any traditional algorithm working on single-layer networks can then be applied. While computationally efficient, this approach discards all information about the distinct layers and their relationships.
Layer-by-layer algorithms first detect communities on each layer independently and then aggregate the results. These are further divided into three classes: frequent subgraph mining Zeng et al. (2006); Berlingerio et al. (2013), feature integration Dong et al. (2013); Nie et al. (2018), and partition integration Cantini et al. (2015a); Tagarelli et al. (2017). Frequent subgraph mining is based on the idea that the nodes of a multilayer community are assigned to the same group in different layers. Thus, it relies on identifying groups of nodes that are consistently clustered together across many layers. Feature integration relies on the principle that communities are made of nodes that are similar to each other, something typically measured by looking at the eigenvectors of the layers. Communities are found by grouping nodes according to their features via, e.g., k-means clustering Forgy (1965); MacQueen (1967). Partition integration directly combines the divisions found on the individual layers. For instance, one could derive a consensus partition, which is the division that is closest to all the input ones, on average Strehl and Ghosh (2002).
Figure 3: Multilayer communities. The continuous (red) contours indicate communities. (a) Intralayer communities group state nodes on the same layer. (b) Members of cross-layer communities can be on both the same and different layers. (c) Pillar communities group all the counterparts of the nodes in other layers in the same community.
@@
Many of these methods are a direct extension of single-layer algorithms to the multilayer case. Modularity maximization was generalized by defining a multilayer null model based on random walks, with intralinks and interlinks playing different roles Mucha et al. (2010). Local exploration methods, like seed set expansion Interdonato et al. (2017), label propagation Boutemine and Bouguessa (2017) or the clique percolation method Afsarmanesh Tehrani and Magnani (2018), as well as random walk-based algorithms such as Infomap De Domenico et al. (2015a) and Walktrap Kuncheva and Montana (2015), have also been adapted. Furthermore, statistical inference methods, primarily based on the stochastic block model (SBM), offer a principled framework to detect communities, infer relationships between layers, and even test whether an observed single-layer network is better explained as an aggregation of hidden layers Peixoto (2015); De Bacco et al. (2017).
Despite this progress, significant challenges remain. First, a persistent issue is the rigorous assessment of statistical significance. Without it, one cannot distinguish meaningful communities from patterns arising by chance. Second, the increasing availability of node and edge metadata is often used only for validating results, a procedure known to be unreliable Hric et al. (2014); Peel et al. (2017). Instead, this information could be directly integrated into the detection process to find more accurate and relevant communities, a task for which SBM-based approaches are particularly well-suited Hric et al. (2016); Newman and Clauset (2016). Finally, the vast majority of methods are designed for multiplex networks. There is a clear need for algorithms that can handle more general multilayer networks (often called heterogeneous networks in computer science) with different node sets and relationships in each layer, a promising direction for future research Sun and Han (2013).
@@ II.2.2 Dynamics and Processes
In some fields, particularly computer science, it is common to distinguish between actors (the underlying real-world entities) and nodes (their representations within specific layers) Czuba et al. (2024); Zhong et al. (2022); Czuba et al. (2025). While this specific terminology may be less explicit in other domains, the underlying principle of defining processes at the level of the entity, rather than solely at the layer-specific node, is widely relevant. This framework allows dynamics to be driven by the actor, integrating states from all its corresponding nodes simultaneously, or by specific layer-level copies. Consequently, it enables complex activation rules: an actor might change state only if a threshold of its nodes are active (reinforcement), or conversely, activity in a single node might drive the entire actor (contagion).
@@
A paradigmatic case in which multiplexity alters the behaviour of a well-understood process is that of reaction-diffusion systems. In multiplex networks, reaction and diffusion terms can be assigned to different layers or coupled across them. For instance, in Kouvaris et al. (2015), activator and inhibitor species are placed on separate layers, where each species diffuses only within its own layer but reacts across layers. Remarkably, the presence of interlayer coupling allows for Turing-like pattern formation even when both species have identical mobility rates—something impossible in single-layer networks. This form of topology-driven instability, further explored in Asllani et al. (2014), highlights how multiplexity can give rise to heterogeneous and robust spatial patterns under conditions where the dynamics of isolated layers remain homogeneous.
@@
In multiplex systems, one well-studied example of interacting dynamics is the coupling between epidemic spreading and awareness diffusion Granell et al. (2013). In this model, the disease propagates through physical contact on one layer, while information about the disease (awareness) spreads via virtual communication on a separate layer. The two processes mutually inhibit each other, giving rise to a point at which the presence of awareness fundamentally alters both the epidemic threshold and the eventual prevalence.
@@
Multiplex networks also provide a natural setting for studying feedback between processes on different layers. In Nicosia et al. (2017), a model is introduced in which neural synchronisation on one layer modulates resource allocation on another, and vice versa. This mutual feedback leads to the spontaneous emergence of explosive synchronisation and highly heterogeneous resource distributions.
@@
Multiplex networks often feature layers with different topologies or dynamics, leading to hybrid systems. In Hizanidis et al. (2016), a modular neural network is modelled with electrical synapses within communities and chemical synapses across them, each type represented as a different layer. The interplay between these types of couplings gives rise to chimera-like states, where coherent and incoherent activity coexist across modules. Such hybrid multiplex models are relevant to understanding how local synchrony coexists with global disorder in brain dynamics.
@@
A similar approach is taken in Sadilek and Thurner (2015), where a two-layer Kuramoto model is derived from physiologically motivated equations for cortical dynamics. The model captures transitions between resting and background brain activity, successfully reproducing empirical features that were inaccessible to simpler models. Again, multiplexity allows for distinct coupling schemes across layers (e.g., electrical vs chemical, or fast vs slow processes) that shape the global phase diagram.
@@
Another work examined how two layers not directly connected can synchronise through the mediation of a central relay layer Leyva et al. (2018). Results showed that synchronisation thresholds are reduced in the multiplex setting and that low-degree nodes play a crucial role, while hubs can be “demultiplexed” without disrupting global coherence. More recently, a three-layer multiplex of non-linear oscillators, in which a central drive layer unidirectionally couples to two response layers, has been investigated Vadakkan and Ambika (2025). Such directional coupling induces synchronised oscillations in the response layers, even with amplification relative to the drive, and varying interlayer strength or introducing time-scale mismatches enables control over synchronisation patterns, amplitudes, and dynamical regimes such as quasi-periodicity and generalized synchronisation.
@@ II.3.1 Temporal Networks
Complex systems, from social groups to interacting neurons, are fundamentally dynamic. While classic network science often represents these systems with static graphs, this is an aggregation that loses all temporal information, such as the sequence, duration, or concurrency of interactions. The recent availability of time-resolved data and computational power has propelled the study of temporal networks, revealing phenomena invisible in a static view Holme and Saramäki (2012); Holme (2015). A temporal network can be represented as a sequence of static snapshots, which naturally fits the multilayer network framework where each layer is a snapshot in time. Besides, it is also possible to have networks that are temporal and multilayer at the same time, in which case the temporal correlations between layers can significantly affect the dynamics of processes unfolding on multilayer networks Starnini et al. (2017).
@@
While early theoretical models added temporality by dynamically changing links, particularly in an adaptive fashion Gross and Blasius (2008), the strong expansion of the field has been mostly data-driven Holme and Saramäki (2012); Holme (2015). The availability of high-resolution temporal data, such as the face-to-face contact network in a museum exhibition depicted in Figure 4 and collected by the SocioPatterns collaboration using wearable sensors Cattuto et al. (2010), provided a strong empirical foundation for the field. This data made it clear that static representations not only hide temporal details but can be misleading, as the properties of a networked system can vary dramatically from one moment to the next Masuda and Lambiotte (2016); Saramäki and Moro (2015).
@@
Figure 4: Temporal networks of face-to-face interactions of people on different days of museum visits. In each day, time evolution is represented as a colour code. The data has been recorded by detecting physical proximity using wearable sensors developed by the SocioPatterns collaboration Isella et al. (2011); Van den Broeck et al. (2012). Reprinted figure with permission from the SocioPatterns collaboration.
@@
This data-centric view is what led researchers to analyse temporal networks as a series of snapshots. This approach revealed that while aggregated properties might converge to the static ones Krings et al. (2012), local measures like a node’s centrality could fluctuate significantly, with nodes shifting from central to peripheral roles over time Braha and Bar-Yam (2006, 2009). However, a true paradigm shift occurred when the focus moved from fluctuating static properties to inherently temporal concepts. Researchers recognized that information flows along temporal paths, which are a set of time-respecting sequences of interactions Kempe et al. (2000); Holme (2005). A path that exists in an aggregated static network may be impossible in a temporal one if the interactions do not occur in the correct chronological order Pan and Saramäki (2011). This insight spurred the redefinition of classical centrality measures for the temporal domain Stehlé et al. (2010) and led to the discovery of new phenomena, such as the burstiness of human interactions, where events are clustered in time rather than occurring randomly Barabasi (2005); Karsai et al. (2018).
@@
But not all temporal interactions are equally informative. Patterns may appear briefly, recur, or vanish entirely. Motifs that are evident in aggregated static views of the network may never exist in real time due to improper chronological alignment Holme (2005). To deal with this complexity, researchers have developed various techniques for filtering and pattern detection. At the local scale, temporal backbones highlight node pairs whose interaction frequencies significantly exceed what would be expected under a temporal null model Kobayashi et al. (2019). Temporal motifs identify frequently recurring, time-ordered subgraphs Kovanen et al. (2013), while temporal cores generalize the static core decomposition to uncover dense substructures active over defined time windows Galimberti et al. (2018). Similarly, the temporal rich-club coefficient measures whether highly connected nodes in the aggregate network tend to form simultaneous, persistent connections Pedreschi et al. (2022).
@@
In parallel, significant effort has gone to identifying the relevant timescales that govern the evolution of temporal networks. Early techniques relied on comparing successive time windows Darst et al. (2016), but this overlooked long-range correlations. More sophisticated methods compute pairwise distances or similarities between network snapshots Masuda and Holme (2019), enabling the clustering of recurring states and the identification of their transitions. Other approaches apply auto-correlation metrics, or Fourier transforms to extract temporal correlations and recurrent interaction patterns at different time scales Lacasa et al. (2022); Andres et al. (2024a).
@@
Representing temporal networks in a simple form for analysis has also been a priority. Two lossless, static-like representations have emerged as especially useful. The supra-adjacency representation encodes node-time pairs as nodes in a multilayer structure, linking them according to the observed interactions Valdano et al. (2015); Sato et al. (2021). The event graph, in contrast, treats interactions themselves as nodes, linking temporally adjacent events involving shared participants Kivelä et al. (2018). This latter framework is especially well-suited to model processes constrained by timing, such as epidemic spreading or transit dynamics Badie-Modiri et al. (2022). Another line of work uses graph embedding techniques to map high-dimensional temporal networks into low-dimensional vector spaces Goyal and Ferrara (2018); Cai et al. (2018). These embeddings preserve structural or temporal similarities and support tasks such as link prediction or influence sets Thongprayoon et al. (2023); Torricelli et al. (2020); Dall’Amico et al. (2024). Embeddings also facilitate data compression and learning-based applications, as discussed further in Section II.3.3.
@@
Models of temporal networks have been developed to generate realistic synthetic networks or to establish null benchmarks. Mechanistic models aim to reproduce observed features or test hypotheses about dynamics Holme and Liljeros (2015), while randomized models offer a controlled way to destroy correlations and assess their significance Gauvin et al. (2022). These models range from memoryless activity-driven frameworks Perra et al. (2012); Starnini and Pastor-Satorras (2013) to non-Markovian extensions capturing burstiness and memory effects Moinet et al. (2015). Additional approaches include stochastic block models, priority queue models, models based on Hawkes or other dynamical processes, and memory network models.
@@
The temporal networks framework also adds another layer of complexity to dynamical processes unfolding on them, since one needs to take into account the coupling between the characteristic timescale of the dynamic process and the temporal network’s one. In epidemiology, the coupling between interactions and disease progression profoundly affects spreading outcomes with important consequences for designing effective public health interventions Masuda and Holme (2017); Machens et al. (2013); Koher et al. (2019); Toroczkai and Guclu (2007); Valdano et al. (2015). In social systems, temporal networks are used to model the spread of information, behaviours, and innovations, often through threshold mechanisms distinct from disease transmission Unicomb et al. (2021); Andres et al. (2024b). Temporal opinion dynamics have been explored to understand consensus and polarization Li and Dankowicz (2019); Chu and Porter (2023), while synchronisation phenomena in biology and engineering are modelled as coupled oscillators interacting over time-varying links Ghosh et al. (2022). Applications in transportation include traffic optimisation and infrastructure design Salama et al. (2022).